# YOLOv11 Optimization for Efficient Resource Utilization

### v2

@@ Abstract
The objective of this research is to optimize the eleventh iteration of You Only Look Once (YOLOv11) by developing size-specific modified versions of the architecture. These modifications involve pruning unnecessary layers and reconfiguring the main architecture of YOLOv11. Each proposed version is tailored to detect objects of specific size ranges, from small to large. To ensure proper model selection based on dataset characteristics, we introduced an object classifier program. This program identifies the most suitable modified version for a given dataset. The proposed models were evaluated on various datasets and compared with the original YOLOv11 and YOLOv8 models. The experimental results highlight significant improvements in computational resource efficiency, with the proposed models maintaining the accuracy of the original YOLOv11. In some cases, the modified versions outperformed the original model regarding detection performance. Furthermore, the proposed models demonstrated reduced model sizes and faster inference times. Models weights and the object size classifier can be found in this repository  [^1] https://github.com/AREEG94FAHAD/yolov11.
Keywords: Object detection, YOLOv11, optimization, image size adaptation, scalable detection.
@@ 1 Introduction
From 2016 to 2024, one-stage object detection methods have gained significantly more popularity and attention from researchers compared to two-stage methods. As illustrated in Figure 1, we analyzed data from Google Scholar, collecting information on the number of research papers referencing one-stage and two-stage object detection during this period.
@@
In this paper, we focus on a popular one-stage algorithm, YOLO, specifically version 11, which was developed in 2024 [33, 19]. The main aim of this study is to propose six modified versions of YOLOv11, optimized for computational resources while maintaining the model’s accuracy compared to the original version. Each version is designed to target specific object sizes. This paper can be considered as the YOLOv11 version of the paper [26].
@@
    Each proposed model has been tested and evaluated on a relevant dataset. The dataset was selected based on an object classification program, and the results were evaluated using standard object detection metrics. The performance of the modified models was compared with the original YOLOv11 model and YOLOv8.
The rest of the paper is organized as follows: In Section II, we provide an overview of the YOLOv11 model and its main components. In Section III, we detail the proposed models. In Section IV, we describe the program used to classify the nature of the dataset and the dataset used to evaluate the models. In Section V, we present the main results and comparisons with recent works. In Section VI, we discuss the limitations of the work and suggest future directions. Finally, in the last section, we conclude by summarizing the key findings.
@@ 3 Modified Versions of YOLOv11
To streamline this process, we used a simple program to analyze and provide detailed information about the object sizes in the dataset  [^2] https://github.com/AREEG94FAHAD/yolov11, as discussed in the dataset section. Using these modified versions of YOLOv11 instead of the original architecture reduces the computational cost and model size while maintaining accuracy in most scenarios.
@@ 3.1 YOLOv11-small
Figure 3: YOLOv11-Small: Architecture Optimized for Detecting Small Objects
@@ 3.2 YOLOv11-medium
Figure 4: YOLOv11-Medium: Architecture Optimized for Detecting Medium Objects
@@ 3.3 YOLOv11-large
Figure 5: YOLOv11-Large: Architecture Optimized for Detecting Large Objects
@@ 3.4 YOLOv11-sm
Figure 6: YOLOv11-sm: Architecture Optimized for Detecting small and medium Objects
@@ 3.5 YOLOv11-ml
Figure 7: YOLOv11-sm: Architecture Optimized for Detecting medium and large Objects
@@ 3.6 YOLOv11-sl
Figure 8: YOLOv11-sl: Architecture Optimized for Detecting small and large Objects
@@ 4.1 Object size classifier
WeedCrop [30, 37] 18,693 15,237 3,363 93 YOLOv11-small
@@
Aerial Airport [7, 12] 11,731 9,008 2,682 41 YOLOv11-sm
@@
Face Detection 620 21 0 599 YOLOv11-sl
@@ 4.2 Dataset used
  • WeedCrop [30, 37]: This dataset is used to detect and classify weeds and crops in the agricultural field. All objects in this dataset are categorized into two classes: weed or crop. It comprises 1,176 images containing approximately 7,853 objects. Augmentations such as rotations, shearing, and brightness adjustments were applied to expand the dataset, increasing the total number of objects to 18,693. According to Table 2, most object instances in this dataset are categorized as small-sized. Therefore, this dataset will be used in the following section to evaluate the YOLOv11-small model.
  • BCCD [29, 21]: The BCCD dataset is the second dataset used for classifying blood cells. It comprises 364 images with approximately 4,888 objects, commonly used in the medical field. The objects belong to three main classes: ’Platelets,’ ’RBC,’ and ’WBC.’ Augmentation techniques, including flips, rotations, cropping, and adjustments to hue, saturation, brightness, and exposure, expanded the dataset to 874 images with a total of 11,780 objects. As shown in Table 2, approximately 86% of all object instances in the BCCD dataset are classified as medium-sized according to Table 1. Therefore, this dataset will be used to evaluate the medium variant of YOLOv11 (i.e., YOLOv11-medium).
@@
  • Brain Tumor [1, 22]: The primary purpose of this dataset is to detect brain tumors. It comprises 9,900 images categorized into three classes. Across all splits—training, testing, and validation—the dataset contains a total of 21,888 object instances. Approximately 95% of the object instances in the Brain Tumor dataset fall within the medium and large size ranges, as shown in Tables 1 and 2. Therefore, the YOLOv11-sm modified version is used for the evaluation process of this dataset.
  • Face Detection [9]: The final dataset used in this paper is the Face Detection dataset, which consists of 389 images, each containing multiple face instances. The model used for this dataset is YOLOv11-sl, designed for both small and large objects, as most object instances in the dataset fall within these two size ranges.
@@ 5.1 Accuracy and Detection Metrics
In this section, we evaluate the model performance using metrics such as Recall, Precision, mAP@50. Three models have been used for the evaluation: the first is the YOLOv11 model, the second is a modified version, and the third is the YOLOv8. Each dataset was tested with each model, resulting in a total of 18 experiments.
Table 4: Performance comparison of YOLOv11, proposed versions, and YOLOv8 across various datasets.
@@
WeedCrop[30, 37] YOLOv11 67.07 70.62 69.09
YOLOv11-small 70.98 69.27 69.27
YOLOv8 63.98 63.98 68.50
@@
As shown in Table 4, in most scenarios, the YOLOv11 and the modified version of YOLOv11 outperforms YOLOv8 across all computed metrics, including Recall, Precision, and mAP@50. However, the performance improvement generally does not exceed 5%. An exception is observed in the BCCD dataset, where YOLOv8 achieves slightly better performance, but the margin is less than 1%. Regarding the performance of the original model (YOLOv11) and the proposed modified models, the results indicate that there is no significant difference in their performance. In four cases, the modified models outperform YOLOv11, as shown in the table. Specifically, YOLOv11-small, YOLOv11-large, YOLOv11-sm, and YOLOv11-sl demonstrate better performance in terms of mAP@50. Conversely, the original YOLOv11 outperforms the modified versions (YOLOv11-medium and YOLOv11-sm). Despite these variations, the maximum performance difference across all metrics is less than 2%.
@@ 5.2 Model Efficiency and Resource Utilization
This section presents the performance of the modified models compared to YOLOv8 and YOLOv11 in terms of model size, GFLOPs, inference time, power consumption.
Figure 10 illustrates the sizes of the models (weights and biases) required for deployment on devices. The modified versions of YOLOv11 demonstrate significant improvements in reducing model size. As shown, all the proposed modified versions are smaller than the original YOLOv11 model. Additionally, the results indicate that YOLOv11, with a size of 5.5, is more compact compared to YOLOv8, which has a size of 6.3.
@@
Regarding GFLOPS, which measures the abbreviation measures of giga floating-point operations per second, the results indicate that YOLOv8 requires higher GFLOPS compared to other models see Figure 11. In contrast, YOLOv11 and the proposed modified versions demonstrate better efficiency. When comparing YOLOv11 to its modified versions, the results highlight a substantial reduction in GFLOPS for the modified versions.
@@
The average inference time highlights the better performance of the proposed modified versions of YOLOv11 compared to the original YOLOv11 and YOLOv8. As shown in Figure 12, the inference time of the modified versions is consistently lower across all scenarios. Additionally, all versions, including the original models, achieved an inference time of less than 5 milliseconds. The difference in inference time across all models for the same dataset is approximately 2 milliseconds.
@@
Figure 13 illustrates the power consumption results for each model, calculated per epoch. As shown, four of the proposed models (YOLOv11-small, YOLOv11-medium, YOLOv11-large, and YOLOv11-ml) outperform the original YOLOv11 and YOLOv8 in terms of power efficiency. Meanwhile, YOLOv11-sm and YOLOv11-sl demonstrate performance comparable to the other models.
@@ 6 Discussion
    Power Usage: The results indicate that reducing the model architecture decreases power consumption during training. As shown, YOLOv11-small, YOLOv11-medium, YOLOv11-large, and YOLOv11-ml achieved lower power consumption compared to both the original YOLOv11 and YOLOv8. Meanwhile, YOLOv11-sm and YOLOv11-sl demonstrated approximately the same power consumption as the original YOLOv11 and YOLOv8.

### v3

@@ Abstract
The objective of this research is to optimize the eleventh iteration of You Only Look Once (YOLOv11) by developing size-specific modified versions of the architecture. These modifications involve pruning specific layers and reconfiguring the main architecture of YOLOv11. Each proposed version is tailored to detect objects of specific size ranges, from small to large. To ensure proper model selection based on dataset characteristics, we introduced an object classifier program. This program identifies the most suitable modified version for a given dataset. The proposed models were evaluated on various datasets and compared with the original YOLOv11, YOLOv10, and YOLOv8 models. The experimental results highlight significant improvements in computational resource efficiency, with the proposed models maintaining the accuracy of the original YOLOv11. In some cases, the modified versions even outperformed the original model in terms of detection performance. Furthermore, the proposed models demonstrated reduced model sizes—each of the six proposed models showed a notable reduction compared to the original. Additionally, the required GFLOPs were reduced from 6.3MB (YOLOv11), 5.7MB (YOLOv10) and 8.1MB (YOLOv8) to just 3.8MB for the large model. All proposed models also achieved faster inference times, significantly reducing the time required to detect objects in images. Models weights and the object size classifier can be found in this repository  [^1] https://github.com/Zarko-ai/yolov11.
Keywords: YOLO, YOLOv11, Computer Vision, Object detection, YOLOv10, YOLOv8, image size adaptation.
@@ 1 Introduction
From 2016 to 2024, one-stage object detection methods have gained significantly more popularity and attention from researchers compared to two-stage methods. As illustrated in Figure 1, We present the number of research papers referencing one-stage and two-stage object detection during this period.
@@
In this paper, we focus on a popular one-stage algorithm, YOLO—specifically version YOLOv11n, developed in 2024 [32, 19]. The primary objective of this study is to propose six modified versions of YOLOv11n, each optimized for computational efficiency while preserving accuracy. Rather than utilizing the full YOLOv11n model, we adapt and tailor each version to detect objects of specific sizes more effectively. Although our proposed modifications are based on the n version, they can also be applied to the other YOLOv11 variants (s, m, x, l). The following are the main contributions of this work.
@@
    Each proposed model has been tested and evaluated on a relevant dataset. The dataset was selected based on an object classification program, and the results were evaluated using standard object detection metrics. The performance of the modified models was compared with the original YOLOv11, YOLOv10, and YOLOv8 Models.
  3. 3.
    The proposed modified versions are applied to YOLOv11n and can also be adapted to other YOLOv11 variants, including YOLOv11s, YOLOv11m, YOLOv11l, and YOLOv11x.
The rest of the paper is organized as follows: In Section 2, we provide an overview of the YOLOv11 model and its main components. In Section 3, we detail the proposed models. In Section 4, we describe the program used to classify the nature of the dataset and the dataset used to evaluate the models. In Section 5, we present the main results and comparisons with recent works. In Section 6, we discuss the limitations of the work and suggest future directions. Finally, in the last section, we conclude by summarizing the key findings.
@@ 3 Modified Versions of YOLOv11
To streamline this process, we used a simple program to analyze and provide detailed information about the object sizes in the dataset  [^2] https://github.com/Zarko-ai/yolov11, as discussed in the dataset section. Using these modified versions of YOLOv11 instead of the original architecture reduces the computational cost and model size while maintaining accuracy in most scenarios.
@@ 3.1 YOLOv11-small
Figure 3: YOLOv11-Small Architecture for Small Object Detection
@@ 3.2 YOLOv11-medium
Figure 4: YOLOv11-Medium Architecture for Medium Object Detection
@@ 3.3 YOLOv11-large
Figure 5: YOLOv11-Large Architecture for Large Object Detection
@@ 3.4 YOLOv11-sm
Figure 6: YOLOv11-sm Architecture for Small and Medium Object Detection
@@ 3.5 YOLOv11-ml
Figure 7: YOLOv11-ml Architecture for Medium and Large Object Detection
@@ 3.6 YOLOv11-sl
Figure 8: YOLOv11-sl Architecture for Small and Large Object Detection
@@ 4.1 Object size classifier
The program is used to analyze the nature of the dataset by computing the size of each object based on Table 1. The size of each object instance is calculated from the label files and then classified according to the criteria defined in Table 1.
@@
WeedCrop [29] 18,693 15,237 3,363 93 YOLOv11-small
@@
Aerial Airport [12] 11,731 9,008 2,682 41 YOLOv11-sm
@@
Face Detection [9] 620 21 0 599 YOLOv11-sl
@@ 4.2 Dataset used
  • WeedCrop [29]: This dataset is used to detect and classify weeds and crops in the agricultural field. All objects in this dataset are categorized into two classes: weed or crop. It comprises 1,176 images containing approximately 7,853 objects. Augmentations such as rotations, shearing, and brightness adjustments were applied to expand the dataset, increasing the total number of objects to 18,693. According to Table 2, most object instances in this dataset are categorized as small-sized. Therefore, this dataset will be used in the following section to evaluate the YOLOv11-small model.
  • BCCD [28, 21]: The BCCD dataset is the second dataset used for classifying blood cells. It comprises 364 images with approximately 4,888 objects, the image size were 416x416, This dataset is commonly used in the medical field. The objects belong to three main classes: ’Platelets,’ ’RBC,’ and ’WBC.’ Augmentation techniques, including flips, rotations, cropping, and adjustments to hue, saturation, brightness, and exposure, expanded the dataset to 874 images with a total of 11,780 objects. As shown in Table 2, approximately 86% of all object instances in the BCCD dataset are classified as medium-sized according to Table 1. Therefore, this dataset will be used to evaluate the medium variant of YOLOv11 (i.e., YOLOv11-medium).
@@
  • Brain Tumor [1, 22]: The primary purpose of this dataset is to detect brain tumors. It comprises 9,900 images categorized into three classes. Across all splits—training, testing, and validation—the dataset contains a total of 21,526 object instances. Approximately 95% of the object instances in the Brain Tumor dataset fall within the medium and large size ranges, as shown in Tables 1 and 2. Therefore, the YOLOv11-sm modified version is used for the evaluation process of this dataset.
  • Face Detection [9]: The final dataset used in this paper is the Face Detection dataset, which consists of 389 images, each containing multiple face instances, with a total of 620 faces. The model used for this dataset is YOLOv11-sl, designed for both small and large objects, as most object instances in the dataset fall within these two size ranges.
@@ 5.1 Accuracy and Detection Metrics
In this section, we evaluate the model performance using metrics such as Recall, Precision, mAP@50. Four models have been used for the evaluation: the first is the YOLOv11 model, the second is a modified version, and the third is the YOLO10 and the last one is YOLOv8. Each dataset was tested with each model, resulting in a total of 24 experiments.
Table 4: Performance comparison of YOLOv11, proposed versions, YOLOv10, and YOLOv8 across various datasets.
@@
WeedCrop[29] YOLOv11 66.94 74.58 72.55
YOLOv11-small 70.56 72.09 73.58
YOLOv10 70.69 70.04 72.22
YOLOv8 67.73 63.42 69.41
@@
As shown in Table 4, in most scenarios, the YOLOv11 and the modified version of YOLOv11 outperforms YOLOv10 and YOLOV8 across all computed metrics, including Recall, Precision, and mAP@50. However, the performance improvement generally does not exceed 5%. An exception is observed in the BCCD dataset, where YOLOv8 achieves slightly better performance, but the margin is less than 1%. Regarding the performance of the original model (YOLOv11) and the proposed modified models, the results indicate that there is no significant difference in their performance. In four cases, the modified models outperform YOLOv11, as shown in the table. Specifically, YOLOv11-small, YOLOv11-large, YOLOv11-sm, and YOLOv11-sl demonstrate better performance in terms of mAP@50. Conversely, the original YOLOv11 outperforms the modified versions (YOLOv11-medium and YOLOv11-sm). Despite these variations, the maximum performance difference across all metrics is less than 2%.
@@ 5.2 Model Efficiency and Resource Utilization
This section presents the performance of the modified models compared to YOLOv11, YOLOv10, and YOLOv8 in terms of model size, GFLOPs, inference time, power consumption.
Figure 10 illustrates the sizes of the models (weights and biases) required for deployment on devices. The modified versions of YOLOv11 demonstrate significant improvements in reducing model size. As shown, all the proposed modified versions are smaller than the original YOLOv11, YOLOv10, and YOLOV8 models.
@@
Regarding GFLOPS, which measures the abbreviation measures of giga floating-point operations per second, the results indicate that YOLOv10 and YOLOv8 requires higher GFLOPS compared to other models see Figure 11. In contrast, YOLOv11 and the proposed modified versions demonstrate better efficiency. When comparing YOLOv11 to its modified versions, the results highlight a substantial reduction in GFLOPS for the modified versions.
@@
The average inference time highlights the better performance of the proposed modified versions of YOLOv11 compared to the original YOLOv11, YOLOv10 and YOLOv8. As shown in Figure 12, the inference time of the modified versions is consistently lower across all scenarios. Additionally, all versions, including the original models, achieved an inference time of less than 5 milliseconds. The difference in inference time across all models for the same dataset is approximately 3 milliseconds.
@@
Figure 13 illustrates the power consumption results for each model, calculated per epoch. As shown, four of the proposed models (YOLOv11-small, YOLOv11-medium, YOLOv11-large, and YOLOv11-ml) outperform the original YOLOv11, YOLOv10, and YOLOv8 in terms of power efficiency. Meanwhile, YOLOv11-sm and YOLOv11-sl demonstrate performance comparable to the other models.
@@ 6 Discussion
    Power Usage: The results indicate that reducing the model architecture decreases power consumption during training, this is due to the reduction in the number of operations required when the full architecture is used. As shown, YOLOv11-small, YOLOv11-medium, YOLOv11-large, and YOLOv11-ml achieved lower power consumption compared to both the original YOLOv11, YOLOv10, and YOLOv8. Meanwhile, YOLOv11-sm and YOLOv11-sl demonstrated approximately the same power consumption as the original YOLOv11 and YOLOv8.
@@ 8 Conclusion
Author‘s contribution Areeg Fahad Rasheed carried out the implementation and writing along with M. Zarkoosh.
Data availability
The datasets mentioned in this paper are available through the cited sources.
Declarations
Conflict of interest
The authors declare that they have no conflict of interest
Ethical approval
Not applicable
Code Availability
The code used for the analysis and experiments in this study will be made available upon request and publication.